OpenStudy (anonymous):

Convert to a common log: @ganeshie8 @gingerbby_

OpenStudy (anonymous):

Convert what?

OpenStudy (anonymous):

\[\log_{3} 4 \]

OpenStudy (anonymous):

\[\log_{3} 4=\frac{ \log4 }{ \log3 }\]would you like to know why?

OpenStudy (anonymous):

i knew it but i wasnt sure if that was the answer

OpenStudy (anonymous):

\[\log_{3} 4 \Rightarrow 3^x = 4 \Rightarrow \log 3^x = \log 4 \Rightarrow x \log 3 = \log 4 \Rightarrow x = \frac{\log 4}{\log 3}\]

OpenStudy (anonymous):

thhaaaakksss

OpenStudy (anonymous):

thaks?